### v1

@@ Abstract
$E_{2}$ $\Delta_{2}$(MeV)$=E_{2}-E_{D\bar{D}^{*}}^{q=0}$26.1(9) -25.4(11) -19.0(7) -18.6(8)
$p^{2}(\mathrm{GeV}^{2})$0.050(2) -0.049(2) -0.037(1) -0.036(1)
$p\cot\delta_{0}(p)$(GeV) -0.154(9)(*) -0.146(10)(*) -0.063(11)(*) -0.066(13)
$(p_{\mathrm{lhc}}^{1\pi})^{2}(\mathrm{GeV}^{2})$0.0135(4) -0.0210(4) -0.0292(3) -0.0400(3)
$a_{0}$ (fm) -1.55(10)(*) -1.50(12)(*) -4.03(91)(*) $\mathbf{-4.0(1.0)}$
@@
$E_{B}$ (MeV) $-9.7^{+2.1}_{-2.2}$ (*) $-9.7^{+1.9}_{-2.0}$ (*) $-1.3^{+0.6}_{-0.8}$ (*) $\mathbf{-1.3^{+0.8}_{-1.0}}$
Bound state and the possible lhc issue.— The values of $(p_{\mathrm{lhc}}^{1\pi})^{2}$ at different $m_{\pi}$’s are also shown in Table 2. It is seen that $p^{2}(E_{2})>(p_{\mathrm{lhc}}^{1\pi})^{2}$ is satisfied only for $m_{\pi}=417(1)$ MeV. So we start the discussion of the near-threshold $D\bar{D}^{*}$ scattering amplitude at this $m_{\pi}$ following the strategy in Ref. Prelovsek and Leskovec (2013); Mohler et al. (2013), since the $p\cot\delta_{0}(p)$ derived at $E_{2}$ and $E_{3}$ from Eq. (4) is valid. The effective range expansion (ERE), $p\cot\delta_{0}(p)=\frac{1}{a_{0}}+\frac{1}{2}r_{0}p^{2}$, at $E_{2}$ and $E_{3}$ gives two equations with two unknowns, namely, the $S$-wave scattering length $a_{0}$ and the effect range $r_{0}$. The solutions are
@@
where the large negative $a_{0}$ signals the existence of a bound state. Taking these values as the approximation of those in the infinite volume, we can solve the pole equation of the scattering amplitude $\mathcal{T}\propto(p\cot\delta_{0}(p)-ip)^{-1}$ to obtain the pole momentum $p_{B}=i\kappa_{B}$ with $\kappa_{B}=49.8_{-18.2}^{+17.0}$ MeV. Consequently, we get the binding energy $E_{B}$,
@@
The Weinberg compositeness criterion states that a near-threshold bound state can be viewed as an admixture of a compact state and a two-hadron channel. The generalized Weinberg relation in Ref. Li et al. (2022), which is more proper for the $D\bar{D}^{*}$ single-channel problem here, implies that the compositeness $X$ computed with ERE should be $X\approx 1$ up to ${O}(p^{2})$ according to the small positive $r_{0}$. Therefore, this bound state is predominantly a $D\bar{D}^{*}$ bound state.
@@
If the same procedure is applied to cases at the other three lighter $m_{\pi}$’s, then similar results can be obtained, as shown in Table 2. The “binding energy” $E_{B}=-9.7_{-2.2}^{+2.1}$ at $m_{\pi}=250(3)$ MeV is also consistent with the value $11(7)$ MeV obtained in Ref. Prelovsek and Leskovec (2013) at $m_{\pi}\approx 266$ MeV. However, the $p^{2}$ of $E_{2}$ for these $m_{\pi}$’s are lower than the corresponding lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}$, such that the $p\cot\delta_{0}(p)$ at $E_{2}$ derived by Eq. (4) is questionable. In Refs. Du et al. (2023); Meng et al. (2023), case studies are performed to the OPE lhc effect on the $DD^{*}$ scattering relevant to $T_{cc}^{+}(3875)$ using the lattice data obtained in Ref. Padmanath and Prelovsek (2022). After considering the OPE, these effective theory studies show that, except for the singular behavior close to $(p_{\mathrm{lhc}}^{1\pi})^{2}$ that $p\cot\delta_{0}(p)$ bends up abruptly when $p^{2}$ approaching $(p_{\mathrm{lhc}}^{1\pi})^{2}$ from above, the overall behavior with respect to $p^{2}$ is similar to that of ERE including the values of $p^{2}$ below $(p_{\mathrm{lhc}}^{1\pi})^{2}$. Figure 3 shows the phase shifts (red points), as well as ERE of $p\cot\delta_{0}(p)$ (green bands) for the four $m_{\pi}$’s. In each panel, the ERE curve intersects with the bound state curve $ip$ (dashed green line). If the $D\bar{D}^{*}$ scattering also has the similar feature to that of the $DD^{*}(I=0)$ scattering when considering the OPE lhc, then it is possible that $p\cot\delta_{0}(p)$ has also a pole near the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}$ (red dashed vertical line) and exhibits a rapid increase there to intersect earlier with the bound state curve $ip$ at $p^{2}>(p_{\mathrm{lhc}}^{1\pi})^{2}$. Therefore, it is very likely that there exists a bound state pole at each $m_{\pi}$ with binding energy $E_{B}$ larger ($|E_{B}|$ smaller) than the value in Table 2.
@@
Possible resonance.— Now we consider the physical implication of the energy level $E_{3}$, which lies between but far from the $D\bar{D}^{*}$ threshold and $E_{D\bar{D}^{*}}^{q=1}$. In Ref. Prelovsek and Leskovec (2013), the $E_{3}$ state is interpreted as the upward shifted $D\bar{D}^{*}$ state with the relative momentum $q=0$ owing to the formation of a bound state below the threshold. The physical consideration behind this is Levinson’s theorem that, when $n_{l}$ bound states are formed in the $l$ partial wave, the phase shift of the $l$-wave phase shift satisfies the relation $\delta_{l}(p=0^{+})=n_{l}\pi$ with the convention $\delta(p=\infty)=0$. However, when there exist $n_{b}$ bare states between the threshold and a specific momentum $p_{\mathrm{max}}$, one should use the generalized Levinson’s theorem $\delta_{l}(0^{+})-\delta_{l}(p_{\mathrm{max}})=(n_{l}-n_{b})\pi$ Vidal and LeTourneux (1992); Li and Wu (2022). For the case at hand with $n_{0}=1$, if the bare $\chi_{c1}(2P)$ state exists theoretically, then the phase shift $\delta_{0}(p)$ will undergo an evolution that it starts from $\pi$ at $p=0^{+}$, falls down when the energy increases, and then return to $\pi$ when the energy passes the $\chi_{c1}(2P)$ mass, as is exactly the case (see the value of the phase shift $\delta_{0}$ at $E_{4}$ in Table 2). It is stressed that the position of $E_{4}$ relative to $E_{D\bar{D}^{*}}^{q=1}$ is crucial for this reasoning. By using the ratio method in Eq. (3), we obtain $E_{4}$ right above $E_{D\bar{D}^{*}}^{q=1}$. Actually, this phase change also hints the existence of a resonance. The Breit-Wigner ansatz for a resonance $(m_{R},\Gamma_{R})$ gives
@@
Using $\delta_{0}(E_{3})$ and $\delta_{0}(E_{4})$ as inputs, the resonance paramters $(m_{R},\Gamma_{R})$ at each $m_{\pi}$ are estimated and are collected in Table 3. Figure 4 illustrates the evolution of $\delta_{0}(E)$ with respect to the center-of-mass energy $E$ of $D\bar{D}^{*}$, where red points are the lattice data, the blue bands are plots of Eq. (7) with parameters in Table 3, and the dashed lines indicate the possible phase change below the resonant energy $E_{R}$.
Figure 4: Fitting of resonance using the Breit-Wigner ansatz. In each panel of the figure the red data points represent the phase shifts corresponding to $E_{3}$ and $E_{4}$ in Table 2 at different $m_{\pi}$. The blue band illustrated the resonance using parameters $(m_{R},\Gamma_{R})$. The blue dashed lines illustrate the schematic evolution of phase shifts starting from $\delta_{0}(0^{+})\sim 180^{\circ}$ such that $\delta_{0}(0^{+})-\delta_{0}(p_{\mathrm{max}})\to 0$ when the $D\bar{D}^{*}$ energy $E(p_{\mathrm{max}})$ passes the mass of the would-be $\chi_{c1}(2P)$.
Table 3: Resonance parameters $(m_{R},\Gamma_{R})$.
@@
Phenomenological studies also indicate that when the coupled channel effect of $D\bar{D}^{*}$ and the quark model $\chi_{c1}(2P)$ is considered, apart from a narrow structure near the $D^{0}\bar{D}^{*0}$ threshold, there may exist a wide resonance below 4.0 GeV with the width ranging from 17 to 80 MeV Kalashnikova (2005); Cincioglu et al. (2016); Giacosa et al. (2019); Deng et al. (2023); Wang et al. (2023). These observations are compatible with the results of this study. Experimentally, $X(3940)$ reported by Belle in 2007  Abe et al. (2007); Pakhlov et al. (2008) has the resonance parameters $(m_{X},\Gamma)\sim(3942(9),37_{-17}^{+27})$ MeV, and decays mainly into $D\bar{D}^{*}$ but not $D\bar{D}$, and therefore favors the assignment of a $1^{++}$ resonance. However, it is puzzling that a similar structure has not been observed by other experiments yet.

### v2

@@ II.1 Finite volume energies
$E_{2}$ $\Delta_{2}$(MeV)$=E_{2}-E_{D\bar{D}^{*}}^{q=0}$--26.1(9) −--25.4(11) −--19.0(7) −--18.6(8)
$p^{2}(\mathrm{GeV}^{2})$--0.050(2) −--0.049(2) −--0.037(1) −--0.036(1)
$p\cot\delta_{0}(p)$(GeV) −--0.154(9)(*) −--0.146(10)(*) −--0.063(11)(*) −--0.066(13)
$(p_{\mathrm{lhc}}^{1\pi})^{2}(\mathrm{GeV}^{2})$--0.0135(4) −--0.0210(4) −--0.0292(3) −--0.0400(3)
($E_{2,3}$) $a_{0}$ (fm) −--1.55(10)(*) −--1.50(12)(*) −--4.03(91)(*) −$\mathbf{-4.0(1.0)}$
@@
$p_{B}$ ($\mathrm{MeV}$) $+i~{}137(9)$(*) $+i~{}137(11)$(*) $+i~{}50(11)$ (*) $\mathbf{+i~{}50(13)}$
$E_{B}$ (MeV) −$-9.7^{+2.1}_{-2.2}$ (*) −$-9.7^{+1.9}_{-2.0}$ (*) −$-1.3^{+0.6}_{-0.8}$ (*) −$\mathbf{-1.3^{+0.8}_{-1.0}}$
$K$-matrix $M$ ($\mathrm{MeV}$) 3948(6)(*) 3981(11)(*) 3960(11)(*) $\mathbf{3979(10)}$
($E_{2,3,4}$) $\gamma$--381(22)(*) −--468(43)(*) −--750(101)(*) −$\mathbf{-677(90)}$
$g(\mathrm{GeV}^{2})$--220(14)(*) −--205(25)(*) −--565(97)(*) −$\mathbf{-532(94)}$
$p_{B}$ ($\mathrm{MeV}$) $+i~{}146(9)$(*) $+i~{}141(10)$(*) $+i~{}57(11)$ (*) $\mathbf{+i~{}58(12)}$
$E_{B}$ ($\mathrm{MeV}$) −11(1)$(1)- 1$(*) −$-10(2)$(*) −$-1.7(7)$(*) −$\mathbf{-1.7(7)}$
II.2 Bound state and the possible lhc issue
The values of $(p_{\mathrm{lhc}}^{1\pi})^{2}$ at different $m_{\pi}$’s are also shown in Table 2. It is seen that $p^{2}(E_{2})>(p_{\mathrm{lhc}}^{1\pi})^{2}$ is satisfied only for $m_{\pi}=417(1)~{}\text{MeV}$. So we discuss the near-threshold $D\bar{D}^{*}$ scattering amplitude at $m_{\pi}=417(1)~{}\text{MeV}$ following the strategy in Ref. Prelovsek and Leskovec (2013); Mohler et al. (2013), since the $p\cot\delta_{0}(p)$ derived at $E_{2}$ and $E_{3}$ from Eq. (3) are valid. The effective range expansion (ERE), $p\cot\delta_{0}(p)=\frac{1}{a_{0}}+\frac{1}{2}r_{0}p^{2}$, at $E_{2}$ and $E_{3}$ gives two equations with two unknowns, namely, the $S$-wave scattering length $a_{0}$ and the effect range $r_{0}$. The solutions are
@@ II.2 Bound state and the possible lhc issue
where the large negative $a_{0}$ signals the existence of a bound state. Taking these values as the approximation of those in the infinite volume, the scattering amplitude ${t}\propto(p\cot\delta_{0}(p)-ip)^{-1}$ implies a bound state pole at $p_{B}=i\kappa_{B}$ with $\kappa_{B}=49.8_{-18.2}^{+17.0}$ MeV. Consequently, we get the binding energy $E_{B}$,
@@
Weinberg’s compositeness criterion states that a near-threshold bound state can be viewed as an admixture of a compact state and a two-hadron channel. According to the generalized Weinberg relation in Ref. Li et al. (2022), the small positive $r_{0}$ indicates the compositeness $X\approx 1$ up to ${O}(p^{2})$. Therefore, this bound state is predominantly a $D\bar{D}^{*}$ bound state.
@@
If the same procedure is applied to cases at other $m_{\pi}$’s, then similar results can be obtained, as shown in Table 2. The “binding energy” $E_{B}=-9.7_{-2.2}^{+2.1}~{}\mathrm{MeV}$ at $m_{\pi}=250(3)$ MeV is also consistent with the value $11(7)$ MeV obtained in Ref. Prelovsek and Leskovec (2013) at $m_{\pi}\approx 266$ MeV. However, the $p^{2}$ of $E_{2}$ for these $m_{\pi}$’s are lower than the corresponding $(p_{\mathrm{lhc}}^{1\pi})^{2}$ and make the derived $p\cot\delta_{0}(p)$ at $E_{2}$ questionable. Refs. Du et al. (2023); Meng et al. (2024) discuss the OPE lhc effect on the $T_{cc}^{+}(3875)$ relevant $DD^{*}$ scattering using the lattice data in Ref. Padmanath and Prelovsek (2022) and find that, except for the singular behavior that $p\cot\delta_{0}(p)$ bends up abruptly when $p^{2}$ approaching $(p_{\mathrm{lhc}}^{1\pi})^{2}$ from above, the overall $p^{2}$ behavior is similar to that of ERE including the values of $p^{2}$ below $(p_{\mathrm{lhc}}^{1\pi})^{2}$. Figure 3 shows the phase shifts (red points), as well as ERE of $p\cot\delta_{0}(p)$ (green bands) for the four $m_{\pi}$’s. In each panel, the ERE curve intersects with the bound state curve $ip$ (dashed green line). If the $D\bar{D}^{*}$ scattering also has the similar feature to that of the $DD^{*}(I=0)$ scattering when considering the OPE lhc, then it is possible that $p\cot\delta_{0}(p)$ has also a pole near the lhc point $(p_{\mathrm{lhc}}^{1\pi})^{2}$ (red dashed vertical line) and exhibits a rapid increase there to intersect earlier with the bound state curve $ip$ at $p^{2}>(p_{\mathrm{lhc}}^{1\pi})^{2}$. Therefore, it is very likely that there exists a bound state pole at each $m_{\pi}$ with a smaller $|E_{B}|$ than the value in Table 2.
Figure 4: The $S$-wave $D\bar{D}^{*}$ scattering phase $\delta_{0}$ at $E_{3,4}$ and the Breit-Wigner parameterization. In each panel of the figure, the red data points represent the phase shifts at $E_{3}$ and $E_{4}$ at each $m_{\pi}$ (see Table 2). The vertical line indicates the $D\bar{D}^{*}$ threshold. The blue bands (labelled by BW($E_{2,3}$)) illustrate the resonance using parameters $(m_{R},\Gamma_{R})$ from $E_{2,3}$ based on the Breit-Wigner ansatz in Eq. (6).
II.3 Possible resonance
@@
Possible resonance.— Now we consider the physical implication of the energy level $E_{3}$, which lies in the middle of the $D\bar{D}^{*}$ threshold and $E_{D\bar{D}^{*}}^{q=1}$. In Ref. Prelovsek and Leskovec (2013), the $E_{3}$ state is interpreted as the upward shifted $D\bar{D}^{*}$ state with the relative momentum $q=0$ owing to the formation of a bound state below the threshold. The physical consideration behind this is Levinson’s theorem that, when $n_{l}$ bound states are formed in the $l$ partial wave, the phase shift satisfies the relation $\delta_{l}(p=0^{+})=n_{l}\pi$ with the convention $\delta(p=\infty)=0$. However, when there exist $n_{b}$ confined channels (with discretized energies) between the threshold and a specific momentum $p_{\mathrm{max}}$, one should use the generalized Levinson’s theorem $\delta_{l}(0^{+})-\delta_{l}(p_{\mathrm{max}})=(n_{l}-n_{b})\pi$ Vidal and LeTourneux (1992); Li and Wu (2022). For the case at hand with $n_{0}=1$ and the possible existence of $\chi_{c1}(2P)$ (the confined channel) expected by quark models, $\delta_{0}(p)$ will undergo an evolution that it starts from $\pi$ at $p=0^{+}$ ($\cot\delta_{0}(0^{+})=\lim_{p\to 0^{+}}1/(a_{0}p)\to-\infty$ for $a_{0}<0$), falls down when the energy increases, and then return to $\pi$ when the energy passes the $\chi_{c1}(2P)$ mass, as is exactly the case (see the values of the phase shift $\delta_{0}\to 180^{\circ}$ at $E_{4}$ in Table 2). It is stressed that $E_{4}>E_{D\bar{D}^{*}}^{q=1}$ is crucial for this reasoning. Actually, this phase change also hints at the existence of a resonance. The Breit-Wigner ansatz for a resonance $(m_{R},\Gamma_{R})$ gives
@@
Using $\delta_{0}(E_{3})$ and $\delta_{0}(E_{4})$ as inputs, the resonance parameters $(m_{R},\Gamma_{R})$ at each $m_{\pi}$ are estimated and are collected in Table 3. Figure 4 illustrates the evolution of $\delta_{0}(E)$ with respect to the center-of-mass energy $E$ of $D\bar{D}^{*}$, where red points are the lattice data, the blue bands are plots of Eq. (6) with parameters in Table 3, and the dashed lines indicate the possible phase change below the resonant energy $E_{R}$. Since they are tentatively estimated from only two lattice energy levels, the resonance parameters $(m_{R},\Gamma_{R})$ may change if more scrutinized lattice investigations are performed.
Phenomenological studies also indicate that when the coupled-channel effect of $D\bar{D}^{*}$ and the quark model $\chi_{c1}(2P)$ is considered, apart from a narrow structure near the $D^{0}\bar{D}^{*0}$ threshold, there may exist a wide resonance below 4.0 GeV with the width ranging from 17 to 80 MeV Kalashnikova (2005); Cincioglu et al. (2016); Giacosa et al. (2019); Deng et al. (2023); Wang et al. (2023, 2024). These observations are compatible with the results of this study. Experimentally, $X(3940)$ reported by Belle in 2007  Abe et al. (2007); Pakhlov et al. (2008) has the resonance parameters $(m_{X},\Gamma)\sim(3942(9),37_{-17}^{+27})$ MeV, and decays mainly into $D\bar{D}^{*}$ but not $D\bar{D}$, and therefore favors the assignment of a $1^{++}$ resonance. However, it is puzzling that a similar structure has not been observed by other experiments yet.
Table 3: Bound state and resonance parameters. The first group of the data lists the resonance parameters $(m_{R},\Gamma_{R})$ from $E_{3,4}$ are obtained by Eq. (6). The second group shows the pole parameters and the pole residuals of the bound state pole through the $K$-matrix paramterization of the phase shifts at $E_{2,3,4}$. The third group gives the resonance parameters as well as the resonance coupling $|c_{0}|^{2}$ to $D\bar{D}^{*}$, the partial decay width $\Gamma_{D\bar{D}^{*}}$ and the corresponding branching ratio.
@@
Bound state pole and residual from $E_{2,3,4}$
$\mathrm{Im}p_{B}$(MeV) $+i$ 146(9) $+i$ 141(10) $+i$ 57(11) $+i$ 58(12)
$E_{B}$(MeV) −--11(1) −--10(2) −--1.6(7) −--1.7(7)
$c_{0}^{2}$(GeV2) 62(4) 58(4) 23(5) 24(5)
$c_{0}$(GeV) 7.9(3) 7.6(3) 4.8(5) 4.9(5)
$\hat{g}$ 2.04(7) 1.97(8) 1.23(12) 1.24(14)
Resonance pole and residue from $E_{2,3,4}$
$\mathrm{Re}p_{R}$(MeV) 545(4) 559(6) 569(4) 563(4)
$\mathrm{Im}p_{R}$(MeV) −$-i$ 56(6) −$-i$ 34(8) −$-i$ 38(7) −$-i$ 45(7)
$m_{R}$(MeV) 4008(4) 4029(4) 4050(3) 4071(3)
$\Gamma_{R}$(MeV) 60(6) 38(9) 43(8) 50(7)
$|c_{0}|^{2}$(GeV2) 47(5) 28(7) 32(6) 38(6)
$|c_{0}|$(GeV) 6.9(4) 5.3(6) 5.6(5) 6.1(5)
$\hat{g}$ 1.71(9) 1.32(17) 1.39(13) 1.51(12)
$\Gamma_{D\bar{D}^{*}}$(MeV) 63(6) 39(9) 44(8) 51(8)
$\mathrm{Br}_{D\bar{D}^{*}}(\%)$$\sim 100$$\sim 100$$\sim 100$$\sim 100$
II.4 Joint analysis on $E_{2,3,4}$
We also perform a joint analysis based on the energy levels $E_{2,3,4}$. The $K$-matrix parameterization is applied to describe the single channel $S$-wave $D\bar{D}^{*}$ scattering amplitude ${t}$ at $E_{2,3,4}$, namely,
$t_{0}(s)=\frac{K(s)}{1-K(s)i\rho(s)},$ (7)
where $s$ is the invariant mass squared of $D\bar{D}^{*}$, $\rho(s)$ is the two-body phase space factor $\rho(s)=p/(8\pi\sqrt{s})$, and $K(s)$ is a real function of $s$, such that the unitary condition
$\mathrm{Im}{t_{0}}^{-1}(s)=-\rho(s)=-\frac{1}{16\pi}\frac{2p}{\sqrt{s}}\Theta(\sqrt{s}-E_{\mathrm{thr.}}),$ (8)
is satisfied for $s>E_{\mathrm{thr.}}$ with $E_{\mathrm{thr.}}=m_{D}+m_{D^{*}}$. In the case of this study, we have the relation
$p\cot\delta_{0}(p(s))=16\pi\frac{\sqrt{s}}{2}K^{-1}(s).$ (9)
As shown in Fig. 5, the phase shift we obtain indicates $p\cot\delta_{0}(p)\approx 0$ near $p^{2}(E_{3})$, so we parameterize $K(s)$ as
$K(s)=\frac{g}{M^{2}-s}+\gamma$ (10)
with $M,g,\gamma$ being free real parameters, where the single pole term is introduced to account for the possible zero of $K^{-1}(s)$. Thus we tentatively solve the equation given by Eqs. (9) and (10) to determine these parameters using the values of $p\cot\delta_{0}(p)$ at $E_{2,3,4}$. The obtained parameters $M,g,\gamma$ at the four values of $m_{\pi}$ are listed in Table 2, where the errors are obtained through the jackknife analysis.
Figure 5: Phase shifts at $E_{2},E_{3},E_{4}$ and the $K$-matrix parameterization. The layout is similar to that of Fig. 3. The green bands illustrate the $K$-matrix expression of $p\cot\delta_{0}(p)$ $K$-matrix in Eq. 9.
The green bands in Fig. 5 illustrate the function form in Eq. (9) with the determined parameters. It is seen that the green bands intersect with the lower branch of the parabola (green dashed line) and indicate the possible existence of a bound state for each $m_{\pi}$. By solving the pole equation $p\cot\delta_{0}(p)=ip$ in the physical sheet, we obtain positive imaginary values of $p_{B}$ and therefore negative binding energies $E_{B}<0$ for the bound states at the four $m_{\pi}$, which are also listed in Table 2. The values of $E_{B}$ are compatible with the results from the $E_{2,3}$ through ERE. The tiny slopes of the green bands near the threshold also indicate small positive effective ranges $r_{0}$ (if ERE is applicable in this region), which also imply the compositeness $X\approx 1$ of the bound states. Let $s_{0}=(m_{D}+m_{D^{*}}+E_{B})^{2}$ be the bound state pole position in $s$-plane, then close to the pole, the scattering amplitude $t$ has the asymptotic behavior
$t_{0}(s\approx s_{0})=\frac{c_{0}^{2}}{s_{0}-s},$ (11)
with the pole residual $c_{0}^{2}$ giving the coupling of the bound state to $D\bar{D}^{*}$. With the parameterization of $K(s)$ in Eq. (10) and performing a contour integral around $s_{0}$, we obtain positive values of $c_{0}^{2}$, which give the couplings
$c_{0}(m_{\pi}\approx 250~{}\mathrm{MeV})$ $=$ $7.9(3)~{}\mathrm{GeV}$
$c_{0}(m_{\pi}\approx 307~{}\mathrm{MeV})$ $=$ $7.6(3)~{}\mathrm{GeV}$
$c_{0}(m_{\pi}\approx 362~{}\mathrm{MeV})$ $=$ $4.8(5)~{}\mathrm{GeV}$
$c_{0}(m_{\pi}\approx 417~{}\mathrm{MeV})$ $=$ $4.9(5)~{}\mathrm{GeV}.$ (12)
By assuming a pure bound state for $X(3872)$ (the compositeness $X\approx 1$), an effective field theory study comes out with the effective coupling Guo et al. (2013)
$c_{0}^{2}\approx\frac{16\pi}{\mu}\sqrt{\frac{2|E_{B}|}{\mu}}\big{(}m_{D}m_{D^{*}}\sqrt{s_{0}}\big{)},$ (13)
with $\mu=(m_{D}+m_{D^{*}})/(m_{D}m_{D^{*}})$ being the reduced mass of the $D\bar{D}^{*}$ system, which gives $c_{0}\approx(10.6_{-0.3}^{+0.3},10.4_{-0.6}^{+0.5},6.7_{-0.8}^{+0.6},6.7_{-0.8}^{+0.6})~{}\mathrm{GeV}$ at the four pion masses $m_{\pi}=(250,307,362,417)~{}\mathrm{MeV}$. These values are qualitatively compatible with those in Eq. (II.4) from the pole residuals. In addition, these bound state poles also pass the sanity check that the $S$-matrix satisfies the asymptotic behavior $S(p\approx p_{B})\approx\frac{-ig_{B}^{2}}{p-p_{B}}$ with $g_{B}^{2}=c_{0}^{2}\mu/\sqrt{s_{0}}>0$ Iritani et al. (2017).
Figure 6: Pole positions in the complex $p$-plane (left) and $\sqrt{s}$-plane (right). Only physical poles are shown. The bound state poles reside on the positive imaginary axis of the $p$-plane (left) and the negative real axis of $\sqrt{s}$-plane (right). The points in the lower half of the $p$-plane (left) and the sight-bottom $\sqrt{s}$-plane illustrate the resonance poles.
With the parameters $M,g,\gamma$, we check if there exist resonance poles by solving the pole equation $p(s)\cot\delta_{0}(p(s))-ip(s)=0$ in the unphysical Riemann sheet with $\mathrm{Im}p(s)<0$. A pair of conjugate resonance poles $\sqrt{s_{0}}=m_{R}\pm i\frac{\Gamma_{R}}{2}$ are obtained for each $m_{\pi}$,
$(m_{R},\Gamma_{R})(m_{\pi}\approx 250~{}\mathrm{MeV})$ $=$ $(4008(4),60(6))~{}\mathrm{MeV}$
$(m_{R},\Gamma_{R})(m_{\pi}\approx 306~{}\mathrm{MeV})$ $=$ $(4029(4),38(9))~{}\mathrm{MeV}$
$(m_{R},\Gamma_{R})(m_{\pi}\approx 362~{}\mathrm{MeV})$ $=$ $(4050(3),43(8))~{}\mathrm{MeV}$
$(m_{R},\Gamma_{R})(m_{\pi}\approx 417~{}\mathrm{MeV})$ $=$ $(4071(3),50(7))~{}\mathrm{MeV},$
whose positions in the complex $p$-plane and $\sqrt{s}$-plane are illustrated in Fig. 6. We also use Eq. (11) to extract the resonance coupling $c_{0}$ at the pole $s\approx s_{0}$ and obtain complex values of $c_{0}^{2}$ at different $m_{\pi}$. Their modulus $|c_{0}|^{2},|c_{0}|$ and the corresponding dimensionless coupling $\hat{g}=|c_{0}|/m_{R}$ are collected in Table 3. With the coupling $|c_{0}|^{2}$ of the resonance $R$ to $D\bar{D}^{*}$, the partial decay width $\Gamma_{D\bar{D}^{*}}$ is predicted using
$\Gamma_{D\bar{D}^{*}}=|c_{0}|^{2}\frac{\rho(m_{R}^{2})}{m_{R}},$ (15)
which are also listed in Table 3. It is seen that, for all the four $m_{\pi}$, we have $\Gamma_{D\bar{D}^{*}}\approx\Gamma_{R}$, which implies that the decay branching fraction $\Gamma_{D\bar{D}^{*}}/\Gamma_{R}$ is almost 100%, as is expected since $D\bar{D}^{*}$ is the only open-charm decay channel permitted by the kinematics and symmetries.
We also plot in Fig. 7 the squared $D\bar{D}^{*}\to D\bar{D}^{*}$ scattering amplitude $\rho^{2}|t_{0}|^{2}$ for $\sqrt{s}>E_{\mathrm{thr.}}$. For each $m_{\pi}$, the zero value of $\rho^{2}|t_{0}|^{2}$ is due to the well-known Castillejo-Dalitz-Dyson (CDD) zero Castillejo et al. (1956) of the amplitude $t_{0}$, which signals the possible existence of an additional component other than the $D\bar{D}^{*}$ system (the quark model $\chi_{c1}(2P)$ for instance) Chew and Frautschi (1961). We notice that a model study on the two-particle scattering shows the presence of the CDD zero when there exists a bare state in the relevant energy region Li and Wu (2022). Theoretically, the vanishing $t_{0}$ at the CDD zero can be attributed to the cancellation of the $D\bar{D}^{*}$ self-interaction and the $D\bar{D}^{*}$-$\chi_{c1}(2P)$ interaction owing to the interplay of hadron and quark degrees of freedom Hyodo et al. (2008); Baru et al. (2010); Hanhart et al. (2011); Kamiya and Hyodo (2017); Guo and Oller (2016); Kang and Oller (2017). The positions $m_{\mathrm{CDD}}$ of the CDD zeros for the four values of $m_{\pi}$ are determined to be
$m_{\mathrm{CDD}}(m_{\pi}\approx 250~{}\mathrm{MeV})$ $=$ $4020(3)~{}\mathrm{MeV}$
$m_{\mathrm{CDD}}(m_{\pi}\approx 306~{}\mathrm{MeV})$ $=$ $4036(3)~{}\mathrm{MeV}$
$m_{\mathrm{CDD}}(m_{\pi}\approx 362~{}\mathrm{MeV})$ $=$ $4054(3)~{}\mathrm{MeV}$
$m_{\mathrm{CDD}}(m_{\pi}\approx 417~{}\mathrm{MeV})$ $=$ $4076(2)~{}\mathrm{MeV},$ (16)
which are very close to the corresponding resonance masses $m_{R}$ and imply a close connection of the resonance with the possible quark model $\chi_{c1}(2P)$.
Figure 7: Amplitudes $\rho^{2}|t_{0}|^{2}$ for the four pion masses. The zero values of $\rho^{2}|t_{0}|^{2}$ signal the presence of the CDD zero Castillejo et al. (1956) due to the existence of additional components (the quark model $\chi_{c1}(2P)$ for example).
The LHCb collaboration recently reported the clear observation of a $1^{++}$ charmoniumlike state $\chi_{c1}(4010)$ with the resonance parameters $(m_{\chi_{c1}},\Gamma_{\chi{c1}}=(4012.5_{-3.9-3.7}^{+3.6+4.1},62.7_{-6.4-6.6}^{+7.0+6.4})~{}\mathrm{MeV}$ Aaij et al. (2024). It is amazing that the resonance parameters we obtain in Eq. (II.4) are in quite good agreement with $\chi_{c1}(4010)$. We are not sure at present whether this coincidence of the lattice QCD result and the experiment is accidental or a glimpse of the truth. After all, we have only three finite volume energy levels $E_{2,3,4}$ at hand, it is not the time to draw a sound conclusion until more scrutinized lattice QCD studies on this topic are performed in the future.
III Summary